OpenStudy (anonymous):

Verify the identity. cos (x+(pi/2)) = -sin x

OpenStudy (campbell_st):

well use the difference of 2 angles cos(A + B) = cos(A)cos(B) - sin(A)sin(B) thats 1 method

OpenStudy (anonymous):

cos (x+(pi/2)= cos x

OpenStudy (campbell_st):

oops sum not difference

OpenStudy (anonymous):

cos=-sinx thats basically what im trying to figure out now

OpenStudy (campbell_st):

well if you make the subsitution you get \[\cos(x + \frac{\pi}{2}) = \cos(x)\cos(\frac{\pi}{2}) - \sin(x) \sin(\frac{\pi}{2})\] whats the value of cos(pi/2)..? whats the value of sin(pi/2) substitute them and seem what happens

OpenStudy (anonymous):

wait i did the identities wrong

OpenStudy (anonymous):

this is so confusing :(

OpenStudy (campbell_st):

an anternative is to graph the curve cos(x) then graph cos(x + pi/2) pi/2 causes a phase shift of pi/2 to the left.... the resulting cruve... is -sin(x)

OpenStudy (anonymous):

ok